OpenStudy (anonymous):

medal A new video game is expected to sell 120 copies the first hour at a local game store. After that, the sales will follow the function s(x) = 15(x − 1) where x is the number of hours. What is the function that shows total sales, including the first hour?

OpenStudy (anonymous):

answer choices s(x) = 15 (x-1) -120 s(x)= 120x-15 s(x) = 120 +15 (x-1) s(x) = 120-15(x-1)

OpenStudy (djjo):

So the function starts with : s(x) = 120 + ( I put a + after 120 because we know that there are following sales after the first hour ). Then, the sales are s(x) = 15 (x-1). So this is the second part of your function. So what's the answer?

OpenStudy (anonymous):

C

OpenStudy (djjo):

That's right. Do you understand now?

OpenStudy (anonymous):

yes but can you help me with another question

OpenStudy (djjo):

Sure

OpenStudy (anonymous):

Neil is analyzing a quadratic function f(x) and a linear function g(x). Will they intersect

OpenStudy (djjo):

Did they give you table?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

g(x) is (-1 0) (-3 1) (-5 2)

OpenStudy (djjo):

If you look at the graph and graph (-1, 0), (-3, 1), or (-5, 2) which one ends up on f(x)?

OpenStudy (anonymous):

answer choices are Yes , at a point with a postive x-coordinate Yes , at a point with a negative x-coordinate Yes, at a point where x is zero No , they will not intersect

OpenStudy (anonymous):

( -3,1)

OpenStudy (djjo):

That's right, so now we know that it does intersect, but where does it intersect? At zero? At a point with a negative x-coordinate? At a point with a positive x-coordinate?

OpenStudy (anonymous):

at a point with a negative x-coordinate

OpenStudy (djjo):

That's right

OpenStudy (anonymous):

i wish i gave u another medal

OpenStudy (djjo):

I don't care about the medals, I just like helping people

OpenStudy (anonymous):

then can i ask u one final question

OpenStudy (djjo):

Of course

OpenStudy (anonymous):

Which graph represents the function f(x) = log10(x − 2)

OpenStudy (anonymous):

never mind the question

OpenStudy (anonymous):

For an angle θ with the point (16, −30) on its terminating side, what is the value of cosine

OpenStudy (djjo):

Or to make it easier, try checking this out: http://openstudy.com/study#/updates/53c05ddce4b05c273e9be461

OpenStudy (anonymous):

answer choices 16 -30 -15/17 8/17

OpenStudy (anonymous):

i guess 8/17

OpenStudy (djjo):

The answer is 8/17, but I am not very familiar with cosines so if you want to understand why that is the answer click the link: http://openstudy.com/study#/updates/53c05ddce4b05c273e9be461